OpenStudy (anonymous):

Given the polynomial function below, find F(-9). F(x) = -x2 - 10x - 2

OpenStudy (alexandervonhumboldt2):

substitute -9 for x and solve

OpenStudy (alexandervonhumboldt2):

f(-9)=-(-9)^2-10*(-9)-2=?

OpenStudy (anonymous):

oh ok i was clueless

OpenStudy (alexandervonhumboldt2):

Hello Welcome to OpenStudy!!! Since you are new here read the tutorial of how to use openstudy: http://openstudy.com/study#/updates/5655f2d1e4b0af360a7cfaa3

OpenStudy (anonymous):

would the answer happen to be -173

OpenStudy (mathstudent55):

\(f(x) = -x^2 - 10x - 2\) \(f(x) = -(-9)^2 - 10(-9) - 2\) \(f(x) = -(81) - (-90) - 2\) \(f(x) = -81 + 90 - 2\) Answer is NOT -173.